OpenStudy (anonymous):

solve each system of equations. 5/x^2-2/y^2+3=0 3/x^2+1/y^2=7

OpenStudy (anonymous):

you should write the equations in a clear form next time. add some parentheses.

OpenStudy (anonymous):

i think you mean \[(5/x^2)-(2/y^2)+3=0 \] \[(3/x^2)+(1/y^2)=7\]

OpenStudy (anonymous):

multiply 2nd equation by 2. then add 1st and 2nd eqns.

OpenStudy (anonymous):

you get (11/x^2)=11, so x=1.

OpenStudy (anonymous):

if x=1, considering 2nd equation, (3/1^2)+(1/y^2)=7, 1/y^2=4, y=1/2.

OpenStudy (anonymous):

thanks